If the mirror separation d = 50 cm changes by 2 nm, what would be the change in relative frequency? If the laser frequency is νL = 4 x 1014s-1, what would be the value of ∆νL=?



1
Expert's answer
2021-01-18T07:40:08-0500

"\Delta f=\frac{c}{2L}=\frac{\left(3\times10^8\frac{m}{s}\right)}{\left(2\times2\times10^{-9}m\right)}=7.5\times10^{16}s^{-1}"

"vL=4\times10^{14}s^{-1}"

"\Delta vL=(7.5\times10^{16}s^{-1})-(4\times10^{14}s^{-1})=7.46x10^{16}"